If a quotient of a number and 15 is added to 1/5, the results is 14/15. Find the number.
Answer by stanbon(75887)

If a quotient of a number and 15 is added to 1/5, the results is 14/15. Find the number.
------
Let the number be "x":
Equation:
(x/15) + (1/5) = (14/15)
-----
Multiply thru by 15 to get:
x + 3 = 14
x = 11
--------
Cheers,
Stan H.
=============